# TASAR: Transfer-based Attack on Skeletal Action Recognition

### v1

@@ A Post-train Bayesian Perspective on Attack
ST-GCN 2s-AGCN MS-G3D CTR-GCN FR-HEAD SFormer ST-GCN 2s-AGCN MS-G3D CTR-GCN FR-HEAD SFormer
ST-GCN I-FGSM 99.26 11.76 8.33 14.22 16.42 15.44 13.23 96.81 8.82 7.10 13.97 16.42 24.75 14.21
MI-FGSM 100.00 17.76 27.20 14.95 26.59 11.76 19.65 99.63 18.75 28.18 15.07 20.22 23.03 21.05
@@
Table 1: The attack success rate(%) of untargeted transfer-based attacks on NTU60 and NTU120. “Ave” was calculated as the average transfer success rate over all target models except for the surrogate. ‘SFormer’ means SkateFormer.
@@
(D) Implementation Details. Our appended model is a simple two-layer fully-connected layer network. Unless specified otherwise, we use $K=3$ and $M=20$ in Eq. 10 for default and explain the reason in the ablation study later. More implementation details can be found in Appendix C.
Evaluation on Normally Trained Models
@@
Evaluation of Untargeted Attack. As shown in in Tab. 1, TASAR significantly surpasses both S-HAR attacks and transfer-based attacks under the black-box settings, while maintaining comparable white-box attack performance. Specially, TASAR achieves the highest average transfer success rate of 35.5% across different models and datasets, surpassing SMART (Wang et al. 2021) (the SOTA S-HAR attack) and MIG (Ma et al. 2023) (the SOTA transfer-based attack) by a large margin of 23.4% and 8.1% respectively. Moreover, TASAR shows consistent transferability across all surrogate models, target models and datasets. These improvements break the common belief that transfer-based attacks in S-HAR suffer from low transferability and highly rely on the chosen surrogate (Lu et al. 2023).
@@
Evaluation of Targeted Attack. In this section, we focus on targeted attacks under the black-box setting. Improving targeted attack transferability on S-HAR is generally more challenging than untargeted attacks. This is primarily due to the significant semantic differences between the randomly selected class and the original one. Attacking a ‘running’ motion to ‘walking’ is generally easier than to ‘drinking’. This is why targeted attacks have lower success rate than untargeted attacks. However, Tab. 2 shows TASAR still outperforms the baseline under most scenarios. Moreover, TASAR can successfully attack the original class to a target with an obvious semantic gap without being detected by humans. The visual examples can be found in Appendix B.
@@
ST-GCN 2s-AGCN MS-G3D CTR-GCN FR-HEAD
ST-GCN MI-FGSM 27.45 3.06 2.32 1.71 1.71 2.20
@@
Evaluation on Ensemble and Defense Models
Evaluation on Ensemble Models. TASAR benefits from the additional model parameters added by the appended Bayesian components. For a fair comparison, we compare it with SOTA ensemble-based methods, i.e., ENS (Dong et al. 2018) and SVRE (Xiong et al. 2022), and the Bayesian Attack (BA) (Li et al. 2023), because they also benefit from the model size. Unlike BA, which re-trains the surrogate into a BNN, we instead append a small Bayesian component for post-training. ENS and SVRE take three models ST-GCN, MS-G3D and DGNN as an ensemble of surrogate models, while BA and TASAR only take MS-G3D as the single substitute architecture. We choose ST-GCN, 2s-AGCN, MS-G3D, CTR-GCN, FR-HEAD as the target models, and evaluate the average white-box attack success rate (WASR), average black-box attack success(BASR) and the number of parameters in Fig. 3. We can clearly see that TASAR (blue line) achieves the best attack performance under both white-box and black-box settings, with an order of magnitude smaller model size. When using MSG3D (12.78M) as the surrogate model, the Bayesian components appended by TASAR only increase 0.012M parameters of the surrogate size, resulting in a memory cost comparable to that of a single surrogate. In contrast, the Bayesian surrogate model used by BA has 15 times more parameters (255.57M) than the single surrogate.
@@
Surrogate Method TRADES BEAT
ST-GCN MS-G3D CTR-GCN STGCN MS-G3D CTR-GCN
ST-GCN MI-FGSM 3.95 3.75 3.54 96.45 22.29 16.45
SMART 2.81 3.13 1.88 80.13 3.34 2.90
TASAR(Ours) 3.92 4.17 2.94 92.19 60.16 39.84
MS-G3D MI-FGSM 3.02 3.02 2.42 36.89 100.00 30.64
SMART 2.50 3.13 3.13 6.69 82.36 4.01
TASAR(Ours) 12.26 10.29 12.25 59.38 100.00 58.59
ST-GCN MI-FGSM 16.05 5.51 8.46 95.83 30.39 16.05
SMART 12.50 5.78 9.06 73.95 4.68 8.28
TASAR(Ours) 12.50 10.22 12.50 97.98 52.34 19.53
MS-G3D MI-FGSM 23.4 7.59 13.11 28.06 97.54 16.54
SMART 19.45 7.42 11.72 26.71 79.68 13.82
TASAR(Ours) 19.79 14.58 17.71 40.63 100.00 32.29
Table 3: The untarget attack success rate (%) against defense models on HDM05 (top) and NTU 60 (bottom). The surrogate models are normally trained without robustness to adversarial examples, whereas the target models are trained using defense methods.
Evaluation on Defense Models. As BEAT shows high robustness against S-HAR white-box attack (Wang et al. 2023), it is also interesting to evaluate its defense performance against black-box attack. We also employ the adversarial training method TRADES (Zhang et al. 2019) as a baseline due to its robustness in S-HAR (Wang et al. 2023). Obviously, in Tab. 3, TASAR still achieves the highest adversarial transferability among the compared methods against defense models, further validating its effectiveness.
Ablation Study
Dual MCMC Sampling. TASAR proposes a new dual MCMC sampling in the post-train Bayesian formulation (Eq. 10). To see its contribution, we conduct an ablation study on the number of appended models ($K$ and $M$ in Eq. 10). As shown in Tab. 4, compared with vanilla Post-train Bayesian strategy ($M$=0), the dual sampling significantly improves the attack performance. Furthermore, although TASAR theoretically requires intensive sampling for inference, in practice, we find a small number of sampling is sufficient ($K=3$ and $M=20$). More sampling will cause extra computation overhead. So we use $K=3$ and $M=20$ by default.
@@
Table 4: Ablation Study on NTU 60 with ST-GCN as the surrogate. $M$ and $K$ are the number dual MCMC sampling in Eq. 10. To isolate the impact of the number of appended models, we employ TASAR without motion gradient. The contribution of the motion gradient will be discussed in the subsequent ablation experiment.
@@
Temporal Motion Gradient. TASAR benefits from the interplay between temporal Motion Gradient (MG) and Bayesian manner. We hence conduct ablation studies(MG/No MG) to show the effects of motion gradient and report the results in Tab. 5. Compared with TASAR without using motion gradient, TASAR with motion gradient consistently improves the attack success rate in both white box and transfer-based attacks, which shows the benefit of integrating the motion gradient into the Bayesian formulation.
Dataset Surrogate Motion Target

### v5

@@ 4.1 RobustBenchHAR Settings
STGCN 2sAGCN MSG3D CTRGCN FRHEAD SFormer STGCN 2sAGCN MSG3D CTRGCN FRHEAD SFormer
STGCN IFGSM 99.26 11.76 8.33 14.22 16.42 15.44 13.23 96.81 8.82 7.10 13.97 16.42 24.75 14.21
MI 100.00 17.76 27.20 14.95 26.59 11.76 19.65 99.63 18.75 28.18 15.07 20.22 23.03 21.05
@@
(D) Implementation Details. Our appended model is a simple two-layer fully-connected layer network. Unless specified otherwise, we use $K=3$ and $M=20$ in Equation 10 for default and explain the reason in the ablation study later. More implementation details can be found in Appendix D.
4.2 Evaluation on Normally Trained Models
@@
Evaluation of Untargeted Attack. As shown in in Table 1, TASAR significantly surpasses both S-HAR attacks and transfer-based attacks under the black-box settings, while maintaining comparable white-box attack performance. Specifically, TASAR achieves the highest average transfer success rate of 35.5% across different models and datasets, surpassing SMART (Wang et al., 2021) (the SOTA S-HAR attack) and MIG (Ma et al., 2023) (the SOTA transfer-based attack) by a large margin of 23.4% and 8.1% respectively. Moreover, TASAR shows consistent transferability across all surrogate models, target models and datasets. These improvements break the common belief that transfer-based attacks in S-HAR suffer from low transferability and highly rely on the chosen surrogate (Lu et al., 2023).
@@
Evaluation of Targeted Attack. In this section, we focus on targeted attacks under the black-box setting. Improving targeted attack transferability on S-HAR is generally more challenging than untargeted attacks. This is primarily due to the significant semantic differences between the randomly selected class and the original one. Attacking a ‘running’ motion to ‘walking’ is generally easier than to ‘drinking’. This is why targeted attacks have lower success rate than untargeted attacks. However, Table 3 shows TASAR still outperforms the baseline under most scenarios. Moreover, TASAR can successfully attack the original class to a target with an obvious semantic gap without being detected by humans. The visual examples can be found in Figure 4.
4.3 Evaluation on Ensemble and Defense Models
Evaluation on Ensemble Models. TASAR benefits from the additional model parameters added by the appended Bayesian components. For a fair comparison, we compare it with SOTA ensemble-based methods, i.e., ENS Dong et al. (2018) and SVRE Xiong et al. (2022), and the Bayesian Attack (BA) Li et al. (2023), because they also benefit from the model size. ENS and SVRE take three models ST-GCN, MS-G3D and DGNN as an ensemble of surrogate models, while BA and TASAR only take MS-G3D as the single substitute architecture. Unlike BA re-training the surrogate into a BNN, TASAR instead appends a small Bayesian component for post-training. We choose ST-GCN, 2s-AGCN, MS-G3D, CTR-GCN, FR-HEAD as the target models, and evaluate the average white-box attack success rate (WASR), average black-box attack success(BASR) and the number of parameters in Figure 4. We can clearly see that TASAR (blue line) achieves the best attack performance under both white-box and black-box settings, with an order of magnitude smaller model size. When using MSG3D (12.78M) as the surrogate model, the Bayesian components appended by TASAR only increase 0.012M parameters of the surrogate size, resulting in a memory cost comparable to that of a single surrogate. In contrast, the Bayesian surrogate model used by BA has 15 times more parameters (255.57M) than the single surrogate.
Since both BA and TASAR are Bayesian-based attacks, we compare the smoothness of their loss landscape in Figure 2. It can be seen that both BA and TASAR exhibit the ability to smoothen the loss landscape, providing empirical evidence for the Bayesian surrogate’s effectiveness in smoothening the loss surface. Further, TASAR and BA achieve the top-2 performance in transfer-based attacks, highlighting the high correlations between loss smoothness and transferability. Compared to BA, TASAR exhibits a significantly flatter loss landscape, aligning with the higher transfer success rate than BA. The key difference between BA and TASAR is that TASAR samples from a smoothed posterior, which shows the benefit of smoothed posterior sampling for improving adversarial transferability.
@@ 4.3 Evaluation on Ensemble and Defense Models
STGCN 2sAGCN MSG3D CTRGCN FRHEAD
STGCN MI 27.45 3.06 2.32 1.71 1.71 2.20
@@
Table 3: The untarget attack success rate (%) against defense models on HDM05 (top) and NTU 60 (bottom).
Surrogate Method TRADES BEAT
STGCN MSG3D CTRGCN STGCN MSG3D CTRGCN
STGCN MI-FGSM 3.95 3.75 3.54 96.45 22.29 16.45
SMART 2.81 3.13 1.88 80.13 3.34 2.90
TASAR 3.92 4.17 2.94 92.19 60.16 39.84
MSG3D MI-FGSM 3.02 3.02 2.42 36.89 100.00 30.64
SMART 2.50 3.13 3.13 6.69 82.36 4.01
TASAR 12.26 10.29 12.25 59.38 100.00 58.59
STGCN MI-FGSM 16.05 5.51 8.46 95.83 30.39 16.05
SMART 12.50 5.78 9.06 73.95 4.68 8.28
TASAR 12.50 10.22 12.50 97.98 52.34 19.53
MSG3D MI-FGSM 23.4 7.59 13.11 28.06 97.54 16.54
SMART 19.45 7.42 11.72 26.71 79.68 13.82
TASAR 19.79 14.58 17.71 40.63 100.00 32.29
@@
Figure 4: The ground truth label ‘Throw’ can be misclassified as ‘Lie down’ on targeted attack by TASAR. The semantic differences between ground truth labels and target labels are large.
Evaluation on Defense Models. As BEAT shows high robustness against S-HAR white-box attack (Wang et al., 2023), it is also interesting to evaluate its defense performance against black-box attack. We also employ the adversarial training method TRADES (Zhang et al., 2019a) as a baseline due to its robustness in S-HAR (Wang et al., 2023). Obviously, in Table 3, TASAR still achieves the highest adversarial transferability among the compared methods against defense models, further validating its effectiveness.
4.4 Ablation Study
Dual MCMC Sampling. TASAR proposes a new dual MCMC sampling in the post-train Bayesian formulation (Equation 10). To see its contribution, we conduct an ablation study on the number of appended models ($K$ and $M$ in Equation 10). To isolate the impact of the number of appended models, we employ TASAR without motion gradient. The contribution of the motion gradient will be discussed in the subsequent ablation experiment. As shown in Section 4.4, compared with vanilla Post-train Bayesian strategy ($M$=0), the dual sampling significantly improves the attack performance. Furthermore, although TASAR theoretically requires intensive sampling for inference, in practice, we find a small number of sampling is sufficient ($K=3$ and $M=20$). More sampling will cause extra computation overhead. So we use $K=3$ and $M=20$ by default.
Table 4: Ablation Study on NTU 60 with ST-GCN as the surrogate. $M$ and $K$ are the number dual MCMC sampling.
@@ 4.4 Ablation Study
Figure 5: The ablation experiments of motion gradient. ‘MG’/‘No MG’ means whether using motion gradient in TASAR.
@@
Temporal Motion Gradient. TASAR benefits from the interplay between temporal Motion Gradient (MG) and Bayesian manner. We hence conduct ablation studies(MG/No MG) to show the effects of motion gradient and report the results in Figure 5. Compared with TASAR without using motion gradient, TASAR with motion gradient consistently improves the attack success rate in both white box and transfer-based attacks, which shows the benefit of integrating the motion gradient into the Bayesian formulation.
4.5 Surrogate Transferability
It is widely believed that transfer-based attacks in S-HAR are highly sensitive to the surrogate choice (Lu et al., 2023; Wang et al., 2023; 2021). In this subsection, we provide a detailed analysis of the factors contributing to this phenomenon. When looking at the results in Table 1 and the visualization of loss landscape in Figure 2 and Appendix C, we note that loss surface smoothness correlates with the adversarial transferability. For example, CTR-GCN, manifesting smoother regions within the loss landscape, demonstrates higher transferability than ST-GCN and STTFormer. STTFormer trained on NTU 120 has a smoother loss surface than ST-GCN (see Appendix C), resulting in higher transferability than ST-GCN. For NTU 60, STTFormer shows a similar loss surface to that of ST-GCN and exhibits comparable transferability. Therefore, we suspect that the loss surface smoothness plays a pivotal role in boosting adversarial transferability for S-HAR, potentially outweighing the significance of gradient-based optimization techniques. Next, two-stream MS-G3D shows the highest transferability. Unlike other surrogates, which solely extract joint information, MS-G3D uses a two-stream ensemble incorporating both joint and bone features, thereby effectively capturing relative joint movements. In conclusion, we suggest that skeletal transfer-based attacks employ smoother two-stream surrogates incorporating both joint and bone information.
5 Conclusion
6 Acknowledgment
This project has received funding from National Natural Science Foundation of China (No. 62302139, No. 62406320), FRFCU-HFUT (JZ2023HGTA0202, JZ2023 HGQA0101).
Appendix A Ethics and Reproducibility Statement
Ethics Statement. TASAR is capable of generating natural-looking adversarial examples in S-HAR that can transfer across different skeletal classifiers. We acknowledge the possibility that TASAR might pose a significant practical threat to the current S-HAR models. However, we believe that in order to build a reliable and robust action recognizer, it is of great necessity to investigate their vulnerability. Therefore, this paper can raise awareness of vulnerability in existing S-HAR models, which greatly outweighs its risk. TASAR can be employed to evaluate the robustness of skeletal classifiers in real-world applications or improve their robustness through adversarial training.
Reproducibility Statement. To ensure the reproducibility of our work, we have included a comprehensive Reproducibility Statement. For the datasets used in our experiments, all the datasets used in this paper are open dataset and are available to the public. We have provided a thorough description of the data processing steps in the supplementary materials. For the novel model and algorithms presented in this work, we have included a link to the downloadable source code and model checkpoint to use our proposed benchmark and build our approach. The source code and model checkpoint can also be found in the supplementary materials. Additionally, all inference details and mathematical deduction can be found in the Appendix B. This Reproducibility Statement is intended to guide readers to the relevant resources that will aid in replicating our work, ensuring transparency and clarity throughout.
Appendix B Inference Details
The detailed inference process for Post-train Dual Bayesian Motion Attack is outlined in Algorithm 1.
Post-train Dual Bayesian Optimization.
The confidence region of the Gaussian posterior in Equation 8 regulated by $\xi$. As $\Delta\theta^{\prime}$ is sampled from a zero-mean isotropic Gaussian distribution, the inner maximization can be solved analytically:
$\Delta\theta_{*}^{\prime}=\lambda_{\xi,\sigma}\nabla_{\theta^{\prime}}L\left(\mathbf{x},y,\theta,\theta^{\prime}\right)/\left\|\nabla_{\theta^{\prime}}L\left(\mathbf{x},y,\theta,\theta^{\prime}\right)\right\|.$ (15)
Then the gradient of $\nabla_{\theta^{\prime}_{k}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)^{T}\Delta\theta^{\prime}$ in Equation 8 becomes $\nabla_{\theta^{\prime}_{k}}L\left(\mathbf{x},y,\theta,\theta^{\prime}_{k}\right)+\mathbf{H}\Delta\theta^{\prime}_{*}$, in which $\mathbf{H}\Delta\theta^{\prime}_{*}$ can be approximately estimated via the finite difference method:
$\mathbf{H}\Delta\theta^{\prime}_{*}\approx\frac{1}{\gamma}\left(\nabla_{\theta^{\prime}_{k}}\frac{1}{K}\sum_{k=1}^{K}L\left(\mathbf{x},y,\theta,\theta^{\prime}_{k}+\gamma\Delta\theta^{\prime}_{*}\right)\right.\left.-\nabla_{\theta^{\prime}_{k}}\frac{1}{K}\sum_{i=1}^{K}L\left(\mathbf{x},y,\theta,\theta^{\prime}_{k}\right)\right),$ (16)
where $\gamma$ is a small positive constant. Therefore, our final optimization objective is:
$\frac{1}{K}\sum_{k=1}^{K}\nabla_{\theta_{k}^{\prime}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)+(1/\gamma)\left(\nabla_{\theta_{k}^{\prime}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}+\gamma\Delta\theta^{\prime}_{*}\right)\right.\left.-\nabla_{\theta_{k}^{\prime}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right)\right).$ (17)
Followed by Wang & Diao (2023), we use Stochastic Gradient Adaptive Hamiltonian Monte Carlo (SGAHMC) (Springenberg et al., 2016) for the post-train dual Bayesian optimization. Our dual Bayesian optimization assume $\theta^{\prime}$ is sampled from Gaussian posterior, which has a presumed isotropic covariance matrix. In practice, we follow the suggestions from Li et al. (2023) to calculate the mean and the covariance from data by using SWAG (Maddox et al., 2019), as SWAG can offer an improved approximation to the posterior over parameters. While the posterior still relies on Gaussian approximation, it specifically incorporates the SWA(Izmailov et al., 2018) solution as its mean, and decomposes the covariance into a low-rank matrix and a diagonal matrix:
$\theta_{k}^{\prime}\sim\mathcal{N}\left(\theta_{k,\mathrm{SWA}}^{\prime},\bm{\Sigma}_{\mathrm{SWAG}}\right),$
$\bm{\Sigma}_{\mathrm{SWAG}}=\frac{1}{2}\left(\bm{\Sigma}_{\text{diag }}+\bm{\Sigma}_{\text{low-rank }}\right),$ (18)
where $\frac{1}{2}(\geq 0)$ can be set to other coefficients, which represent the scaling factor of SWAG for disassociating the learning rate of the covariance. Note that the posterior discussed in the preceding section is formulated based on the worst cases, thus facilitating its effortless integration with SWAG to expand diversity and flexibility. Since $\theta_{k,\mathrm{SWA}}$ is unknown before training terminates, the dispersion of $\theta_{k}^{\prime}$ in the final Bayesian model originates from the combination of two distinct independent Gaussian distributions, with their covariance matrices aggregated. Consequently Equation 18 becomes:
$\theta_{k}^{\prime}\sim\mathcal{N}\left(\theta_{k,\mathrm{SWA}}^{\prime},\bm{\Sigma}_{\theta_{k}^{\prime}}\right),$
$\bm{\Sigma}_{\theta_{k}^{\prime}}=\alpha\left(\bm{\Sigma}_{\text{diag }}+\bm{\Sigma}_{\text{low-rank }}\right)+\beta\mathbf{I},$ (19)
where $\beta$ controls the covariance matrix of the isotropic Gaussian distribution mentioned before.
1 Input: $\mathbf{x}$: training data; $N_{post-train}$: the number of post-train iterations; $N_{dual}$: the number of Post-train Dual Bayesian optimization iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta_{k}^{\prime}$; $c$: moment update frequency; $K$:the number of appended models;$I$:attack iterations; $\theta$: pre-trained surrogate weights;
2Output: $\left\{\theta_{1}^{\prime}+\Delta\theta_{11}^{\prime},\ldots,\theta_{K}^{\prime}+\Delta\theta_{KM}^{\prime}\right\}$: appended surrogate weights; $\tilde{\mathbf{x}}$: adversarial examples;
// Post-train Bayesian
3 Init: randomly initialize $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
4 for i = 1 to $N_{post-train}$ do
5       for k = 1 to $K$ do
6             Randomly sample a mini-batch data $\{\mathbf{x},y\}$;
7             $\theta_{k_{i}}^{\prime}\leftarrow\theta_{k_{i-1}}^{\prime}-\eta\nabla_{\theta_{k_{i}}^{\prime}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right))$;
8             for t = 1 to $M_{\theta^{\prime}}$ do
9                   Update $\theta_{k}^{\prime}$ via SGAHMC;
10
11             end for
12
13       end for
14
15 end for
// Post-train Dual Bayesian Optimization
16 each $\overline{\theta_{k}^{\prime}}\leftarrow\theta_{k_{0}}^{\prime},\overline{{\theta_{k}^{\prime}}^{2}}\leftarrow{\theta_{k_{0}}^{\prime}}^{2}$;
17 for i = 1 to $N_{dual}$ do
18       for k = 1 to $K$ do
19             Randomly sample a mini-batch data $\{\mathbf{x},y\}$;
20             $\theta_{k_{i}}^{\prime}\leftarrow\theta_{k_{i-1}}^{\prime}-\eta\nabla_{\theta_{k_{i}}^{\prime}}L\left(\mathbf{x},y,\theta,\theta_{k}^{\prime}\right))$;
21             Compute $\Delta\theta_{*}^{\prime}$ via Equation 15;
22             Solving outer min optimization in Equation 8 via Equation 17;
23             if $\operatorname{MOD}(i,c)=0$ then
24                   $n\leftarrow i/c$ ;
25                   $\overline{\theta_{k}^{\prime}}\leftarrow\frac{n\overline{\theta_{k}^{\prime}}+\theta_{k_{i}}^{\prime}}{n+1},\overline{{\theta_{k}^{\prime}}^{2}}\leftarrow\frac{n\overline{{\theta_{k}^{\prime}}^{2}}+{\theta_{k_{i}}^{\prime}}^{2}}{n+1}$;
26
27             end if
28
29       end for
30
31 end for
[…]
33
// Attack
34 Initialization:$\tilde{\mathbf{x}}^{0}=\mathbf{x}$;
35 obtain the time-varying parameters with TV-AR;
36 for i = 1 to I-1 do
37       models=[] for k = 1 to $K$ do
38             for $m=1$ to $M$ do
39                   Draw $\theta_{k}^{\prime}+\Delta\theta_{km}^{\prime}$ in Equation 19;
40                   models.append($\theta_{k}^{\prime}+\Delta\theta_{km}^{\prime}$);
41             end for
42
43       end for
44      Obtain the ensemble gradient;
45       Calculate the motion gradient $\left(\frac{\partial L\left(\tilde{\mathbf{x}}^{i}\right)}{\partial\tilde{x}_{t-1}^{i}}\right)_{d1}$ and $\left(\frac{\partial L\left(\tilde{\mathbf{x}}^{i}\right)}{\partial\tilde{x}_{t-2}^{i}}\right)_{d2}$ with Equation 13 and Equation 14;
46       Update $\tilde{\mathbf{x}}^{i+1}$ via Equation 10;
47
48 end for
49return $\left\{\theta_{1}^{\prime}+\Delta\theta_{11}^{\prime},\ldots,\theta_{K}^{\prime}+\Delta\theta_{KM}^{\prime}\right\}$,$\tilde{\mathbf{x}}$;
Algorithm 1 Inference for Post-train Dual Bayesian Motion Attack
Appendix C Additional Experimental Results and Analysis
C.1 Additional Landscape Visualizations
To explore the sensitivity of transfer-based attacks to surrogate models (Lu et al., 2023; Wang et al., 2023; 2021), we present in Figure 6 and Figure 7 the loss landscapes of different surrogate models trained on NTU60 and NTU120. It is evident that both the post-train Bayesian optimization(PB) and the improved post-train Dual Bayesian optimization(P-DB) can smooth the loss landscape; Notably, surrogates refined by P-DB display smoother loss landscapes, leading to superior transferability over normally trained models and those optimized by PB.
Figure 6: Loss landscapes of trained models with different methods on NTU60. The loss landscape in each plot are generated from the same original data randomly selected from the test dataset of NTU60.PB means the post-train Bayesian optimization, P-DB means the improved post-train Dual Bayesian optimization. The first row, second row and third row represent the loss surface of CTR-GCN, ST-GCN and STTFormer trained normally, as well as PB and P-DB, respectively, with the $z$ axis ranging from 0 to 16.
Figure 7: Loss landscapes of trained models with different methods on NTU120.The loss landscape in each plot are generated from the same original data randomly selected from the test dataset of NTU120. PB means the post-train Bayesian optimization, P-DB means the improved post-train Dual Bayesian optimization.The first row, second row, and third row correspond to the loss surfaces of CTR-GCN, ST-GCN, and STTFormer under normal training, PB and P-DB, respectively. For the normally trained plots, the $z$ axis ranges from 0 to 30, while for PB and P-DB, the range is from 0 to 20.
C.2 Additional Results
The Sensitivity to $\xi$.
In our Post-train Dual Bayesian optimization, we consider the worst-case parameters from the posterior. The confidence region of the Gaussian posterior is regulated by $p(\Delta\theta^{\prime})\geq\xi$, influencing the extent of exploration within the posterior distribution. Therefore, we investigate the relationship between the sensitivity of $\xi$ and the performance of TASAR. Based on our assumption of an isotropic Gaussian distribution, we got the analytical solution of the worst case as below:
$\Delta\theta_{*}^{\prime}=\lambda_{\xi,\sigma}\nabla_{\theta^{\prime}}L\left(\mathbf{x},y,\theta,\theta^{\prime}\right)/\left\|\nabla_{\theta^{\prime}}L\left(\mathbf{x},y,\theta,\theta^{\prime}\right)\right\|.$ (20)
Figure 8: The ground truth label ‘Walk’ can be misclassified as ’Cartwheel’ on targeted attack by TASAR. The semantic differences between ground truth labels and target labels are large.
In Equation 20, $\xi$ can be reparameterized as a hyper-parameter $\lambda_{\xi,\sigma}$. Consequently, we conduct ablation studies to investigate the relationship between the performance of our method and the sensitivity of the hyper-parameter $\lambda_{\xi,\sigma}$. We varied with $\lambda_{\varepsilon,\sigma}\in\{0.001,0.01,0.05,0.1,1,1.5,2\}$ on NTU120 dataset and show the success rates of attacking victim models. We choose ST-GCN as the pre-trained model and the results are shown in Table 5. We found that setting a small value of $\lambda_{\varepsilon,\sigma}$ achieves the best adversarial transferability while maintaining a high benign accuracy. Hence we sample a collection of new surrogates near to the original surrogates and set $\lambda_{\varepsilon,\sigma}=0.001$ as default.
Table 5: Comparsions attack success rate(%) with different $\lambda_{\xi,\sigma}$. The surrogate model is uniformly selected as ST-GCN on NTU120.
$\lambda_{\xi,\sigma}$ Target Accuracy